# Denoising Refinement Diffusion Models for Simultaneous Generation of Multi-scale Mobile Network Traffic

### v1

@@ Abstract
Multi-layer mobile network traffic generation is a key approach to capturing multi-scale network dynamics, supporting network planning, and promoting generative management of mobile data. Existing methods focus on generating network traffic with a single spatiotemporal resolution, making it difficult to achieve joint generation of multi-scale traffic. In this paper, we propose ZoomDiff, a diffusion-based multi-scale mobile traffic generation model. ZoomDiff maps the urban environmental context into network traffic with multiple spatiotemporal resolutions through custom-designed Denoising Refinement Diffusion Models (DRDM). DRDM employs a multi-stage noise-adding and denoising process, enabling different stages to generate traffic with distinct spatial and temporal resolutions. It aligns the progressive denoising process of diffusion models with hierarchical network layers, including BSs, cells, and grids with different granularities. Evaluations on real-world mobile traffic datasets demonstrate that ZoomDiff achieves a performance improvement of at least 18.4% over state-of-the-art baselines on generation tasks at multi-scale traffic. The efficiency and generalization ability are also demonstrated, which indicates that ZoomDiff holds strong potential for generative mobile data management. The code of ZoomDiff is available at https://anonymous.4open.science/r/ZoomDiff-105E/.
@@ I Introduction
Mobile network traffic serves as a vital indicator of both user behavior and network performance [1]. Accurate modeling and understanding of traffic dynamics not only reflect underlying communication activities but also inform essential decisions in network planning and optimization. However, the privacy of communication data and the imperfection of mobile traffic data management makes it difficult to obtain accurate and large-scale traffic datasets. Generative approaches can supplement scarce real-world measurements, enhance simulation fidelity, and accelerate the design and testing of network control strategies in a cost-effective and privacy-preserving manner [2]. This supports a variety of downstream optimization tasks such as Base Station (BS) deployment [3, 4, 5, 6], resource management [7, 8], traffic load balancing [9, 10], and energy-efficient scheduling [11].
@@
Over the past decade, extensive efforts have been devoted to generating mobile traffic data across various spatial scales, ranging from city-wide aggregation to regional and BS level granularity [12, 13, 14, 15]. Such work typically leverages historical traffic patterns and statistical modeling approaches to simulate spatiotemporal traffic characteristics at diverse granularities. These studies lay foundational groundwork by adequately addressing macroscopic traffic generation demands and demonstrating their value in multiple practical network scenarios [16, 17, 18].
@@
However, with the escalating complexity of networks and the emergence of novel mobile services such as immersive XR, autonomous driving, and tactile Internet [19, 20, 21], the demand for mobile traffic data generation exhibits two major trends. First, multi-scale traffic generation. The typical structure of a cellular mobile network, as shown in Fig. 1, consists of multiple layers of network nodes, including BSs, cells, grids, and users. As network structures become more complex, joint optimization across multiple network layers can yield greater performance gains compared with traditional single-layer strategies. This trend requires the support of synthetic traffic across multiple network layers with different spatiotemporal resolutions, while the intrinsic spatiotemporal correlations among these layers impose stringent requirements on the cross-layer organization and management of such data. Second, high-resolution traffic generation. High-resolution traffic patterns at fine spatial granularities, such as grid-level data, are increasingly critical for precision-based network optimization, including micro-level capacity planning, fine-grained coverage optimization, and user-centric network configuration [22, 23]. Yet traffic at refined scales often exhibits intricate spatiotemporal dependencies and distributional complexities, which require generation-based data management systems to effectively capture and model fine-grained spatial correlations and rapid temporal variations in network dynamics.
@@
Unfortunately, the aforementioned research can only generate coarse-grained network traffic with the same single spatiotemporal resolution as the training set. When generating and updating multi-scale traffic data, different models must be used to simulate different network layers, which greatly increases computational overhead. Moreover, generating the complex spatiotemporal distribution of fine-grained traffic exceeds the capabilities of existing methods. It requires advanced user localization techniques [24] or specialized network probes [25], incurring substantial communication overhead and computational costs [26, 27], making it infeasible for continuous collection in large-scale live networks. A promising yet largely unexplored alternative is to design a unified model capable of jointly generating multi-scale mobile traffic, allowing the easily estimated coarse spatiotemporal granularity traffic to progressively guide the generation of finer-grained traffic, thereby achieving multi-scale traffic generation while improving the accuracy of fine-grained traffic synthesis.
@@
Recently, the emergence of diffusion models has offered promising insights for fine-grained traffic generation. Diffusion models operate by first progressively injecting noise into the original data and subsequently denoising step-by-step to reconstruct high-resolution images from pure noise. This denoising paradigm naturally aligns with the goal of fine-grained mobile traffic synthesis. Inspired by this, we aim to explore a control mechanism for the intermediate states in the noise-adding and denoising processes, as shown in Fig. 2, aligning the variation of noise intensity with the change in spatiotemporal resolution of network traffic. We seek to make the denoising process progressively zooms in on the network traffic, revealing increasingly finer spatiotemporal details at each step.
@@
Figure 2: Align the spatiotemporal refinement relationships of multi-scale network traffic with the denoising process of diffusion models.
@@
However, achieving this goal presents two main challenges. First, it remains unclear how to design effective denoising procedures within the diffusion framework to accurately represent multi-scale network traffic. Second, it is challenging to model and preserve the cross-resolution dependencies and statistical constraints between traffic at different granularities.
@@
To address the above challenges, we propose ZoomDiff, a multi-resolution mobile traffic generation model. It achieves a mapping from Urban Environmental Context (UEC) to multi-level network traffic, allowing for a zoom-like progressive generation for traffic with different scales. First, we design Denoising Refinement Diffusion Models (DRDM). It is a multi-stage diffusion framework, enabling controllable denoising intermediates through a Hierarchical Noise-adding Process (HNaP) and a Resolution Refinement Denoising Process (RRDP). Multi-scale traffic with different resolutions can be directly extracted from intermediate states at corresponding diffusion steps. Second, we design a diffusion prior noise mechanism to establish correlations between adjacent noise-adding and denoising stages. This design enables the coarse-resolution traffic to serve as guidance to control the generation of high-resolution traffic. Moreover, we discussed the noise scheduling strategies and the Resolution Greediness Preference of the multi-stage model required for implementing DRDM, and revised the corresponding strategies in conventional diffusion models. We found that the combination of discrete noise intensity, discrete noise-adding, and discrete denoising strategies enables ZoomDiff to achieve optimal multi-scale generation performance. The main contributions can be summarized as follows:
@@
  • To the best of our knowledge, we propose the first unified model capable of simultaneous generation of multi-scale mobile traffic. Leveraging the sophisticated design of DRDM, it aligns the progressive refinement of spatiotemporal scales of network traffic with the denoising process. In other words, multiple intermediate states of the denoising process are constrained to correspond to multi-scale network traffic.
@@
ZoomDiff provides an innovative approach to enhancing multi-scale mobile data management through generative Artificial Intelligence (AI). We further validate that ZoomDiff has excellent performance in the management of traffic refinement and cross-city generalization. This will help construct an effective data platform for joint planning of multi-level mobile networks.
@@ II Related Work
II-A Network Traffic Generation
@@
Mobile traffic generation aims to learn the environment-aware spatiotemporal dynamics by mining the correlation between network traffic and external conditions. Traditional generative algorithms, including mathematical statistics [28, 29] and simulation-based approaches [30, 31, 32], have limited generalization capabilities. With the development of Machine Learning (ML), generative models such as Generative Adversarial Networks (GANs) [12, 33, 34, 35] and Transformers [36, 37] can capture the complex correlations between network demand and environmental features in a data-driven manner. Pandey et al. [33] and Li et al. [34] employed GANs to autoregressively model the temporal dynamics of traffic sequences for synthetic traffic generation. Hui et al. [35] integrated urban environments into Urban Knowledge Graphs (UKGs) for generation guidance. As the stable generative capabilities of diffusion models are validated, recent works have attempted to replace GANs with diffusion-based generators. Zhou et al. [38] applied UKG to the diffusion denoising process for accurate urban population generation. Sivaroopan et al. [39] and Jiang et al. [40] converted one-dimensional network traffic into two-dimensional images to leverage the strong image generation capabilities of diffusion models.
@@
However, existing research primarily focuses on traffic generation at a single spatiotemporal resolution. For the task of generating multi-scale traffic, multiple models or separate generations are typically required. In addition, these models are designed for coarse-scale traffic such as BS or cell levels, and perform poorly when generating fine-grained traffic at small grid scales. In contrast, ZoomDiff can generate multi-scale mobile traffic in a single process and progressively guide the generation of fine-grained traffic through coarse-scale traffic, achieving both improved generation accuracy and enhanced efficiency.
II-B Generative Management of Mobile Data
Generative management of mobile data leverages advances in generative AI to synthesize, augment, and anonymize datasets such as network traffic [41, 42, 43] and mobile trajectories [44, 45, 46, 47, 48], addressing issues of limited data availability and privacy concerns [49]. Traditional simulation [50, 51] or rule-based models [48] often struggled to capture the complex spatiotemporal patterns inherent in mobile data, whereas modern deep generative models such as GANs and diffusion models learn these patterns directly from real data, producing high-fidelity synthetic data that preserve key statistical properties of the original. Thirumuruganathan et al. [52] trained deep generative models to learn the underlying data distribution with high fidelity for approximate query processing in databases. Li et al. [41] proposed IoTGemini, a two-module framework with a Packet-Sequence GAN to generate realistic IoT network traffic. In the realm of human mobility, Zhang et al. [44] use a clustering-based sequence GAN to capture diverse movement modalities and generate synthetic user trajectories that closely emulate real-world travel patterns. These generative AI methods enable scalable data augmentation and privacy-preserving data sharing in mobile computing. [49] highlights that generative approaches not only enhance the realism of synthetic mobile data but also outperform conventional methods in retaining essential spatiotemporal characteristics, thereby significantly advancing mobile data management in terms of fidelity and utility.
@@ III PRELIMINARY
III-A Correlations among Multi-scale Network Traffic
The network traffic across different scales, including the BS traffic, cell traffic, and grid-level traffic, exhibits a top-down aggregation relationship due to the inherent hierarchy among their service areas. Figure 4 shows the temporal decomposition of hourly BS traffic, cell traffic, and 50m grid traffic from three kinds of Area of Interest (AoI) over one week. The trend component reflects the distribution of daily average traffic over seven days, while the seasonal component captures the temporal patterns of daily traffic after removing the trend. To be quantitative, Fig. 3 computes the RV-coefficients [53] between adjacent levels of network traffic. It can be observed that traffic at neighboring scales within the same AOI exhibits high similarity. This provides the theoretical foundation for jointly generating multi-scale traffic, where coarse-grained traffic can guide the generation of fine-grained traffic. The RV-coefficient is defined as:
@@ III-B Problem Formulation
Mobile network traffic at different layers of network nodes reflects user demand distributions at different spatiotemporal scales. Given the UEC $\boldsymbol{c}$, the objective of the model is to learn a mapping from UEC to the corresponding regional mobile traffic $\boldsymbol{x}$, and simultaneously generate multi-level traffic within the region at different spatiotemporal resolutions. Consider a mobile network with $K\geq 2$ levels, let $\boldsymbol{x}_{s_{\delta_{k}}t_{\tau_{k}}}$ denote the mobile traffic at level $k$ with spatial resolution $s_{\delta_{k}}$ and temporal resolution $t_{\tau_{k}}$, where $1\leq k\leq K$, $\delta_{k}\geq\delta_{k-1}$ and $\tau_{k}\geq\tau_{k-1}$. This mapping can be expressed as $\boldsymbol{c}\mapsto(\boldsymbol{x}_{s_{\delta_{1}}t_{\tau_{1}}},...,\boldsymbol{x}_{s_{\delta_{K-1}}t_{\tau_{K-1}}},\boldsymbol{x}_{s_{\delta_{K}}t_{\tau_{K}}})^{T\times H\times W}$, where $T$ denotes the temporal length of the finest traffic series, and $H$, $W$ represent the finest spatial dimensions. As the spatiotemporal resolution increases, the traffic becomes increasingly influenced by fine-grained urban environments and user behavioral characteristics. Therefore, generating high-dimensional spatiotemporal traffic at fine resolutions is a complex task.

### v2

# Denoising Refinement Diffusion Models for Simultaneous Generation of Multi-scale Mobile Traffic

@@ Abstract
The planning, management, and resource scheduling of cellular mobile networks require joint estimation of mobile traffic across different layers and nodes. Mobile traffic generation can proactively anticipate user demands and capture the dynamics of network load. However, existing methods mainly focus on generating traffic at a single spatiotemporal resolution, making it difficult to jointly model multi-scale traffic patterns. In this paper, we propose ZoomDiff, a diffusion-based model for multi-scale mobile traffic generation. ZoomDiff maps urban environmental context into mobile traffic with multiple spatial and temporal resolutions through a set of customized Denoising Refinement Diffusion Models (DRDM). DRDM employs a multi-stage noise-adding and denoising mechanism, enabling different stages to generate traffic at distinct spatiotemporal resolutions. This design aligns the progressive denoising process with hierarchical network layers, including base stations, cells, and grids of varying granularities. Experiments on real-world mobile traffic datasets show that ZoomDiff achieves at least an 18.4% improvement over state-of-the-art baselines in multi-scale traffic generation tasks. Moreover, ZoomDiff demonstrates strong efficiency and cross-city generalization, highlighting its potential as a powerful generative framework for modeling multi-scale mobile network dynamics.
@@ I Introduction
Mobile traffic serves as a vital indicator of both user behavior and network performance [10646623]. Accurate modeling and understanding of traffic dynamics not only reflect underlying communication activities but also enable informed decisions in network planning and optimization. However, the privacy of communication data and the limited availability of large-scale, high-quality traffic measurements make it difficult to obtain comprehensive datasets for analysis and modeling. Generative approaches offer a promising alternative by supplementing scarce real-world measurements, improving simulation fidelity, and facilitating the design and evaluation of network control strategies in a cost-effective and privacy-preserving manner [10.1145/3419394.3423643]. Such generative modeling supports various downstream optimization tasks, including Base Station (BS) deployment [10086045, sleeping-1, sleeping-2, sleeping-3], resource allocation [10224322, 9725256], traffic load balancing [10976420, 10904019], and energy-efficient scheduling [11062900].
@@
Over the past decade, extensive efforts have been devoted to generating mobile traffic data across various spatial scales, ranging from city-wide aggregation to regional and BS level granularity [10.1145/3485983.3494844, 10.1145/3637528.3671544, 10.1145/3703447, 9625773]. Such work typically leverages historical traffic patterns and statistical modeling approaches to simulate spatiotemporal traffic characteristics at diverse granularities. These studies lay foundational groundwork by adequately addressing macroscopic traffic generation demands and demonstrating their value in multiple practical network scenarios [9844048, 10234717, 10201864].
@@
However, with the escalating complexity of networks and the emergence of novel mobile services such as immersive XR, autonomous driving, and tactile Internet [10.1145/3636534.3697422, 10466554, 8584062], the demand for mobile traffic data generation exhibits two major trends. First, multi-scale traffic generation. The typical structure of a cellular mobile network, as shown in Fig. 1, consists of multiple layers of network nodes, including BSs, cells, grids, and users. As network structures become more complex, joint optimization across multiple network layers can yield greater performance gains compared with traditional single-layer strategies. This trend requires the support of synthetic mobile traffic across multiple network layers. The spatiotemporal resolution differences and inherent correlations among traffic at these layers impose stringent requirements on the cross-layer joint generation and organization of such traffic data. Second, high-resolution traffic generation. High-resolution traffic patterns at fine spatial granularities, such as grid-level data, are increasingly critical for precision-based network optimization, including micro-level capacity planning, fine-grained coverage optimization, and user-centric network configuration [5970249, 10198898]. Yet traffic at refined scales often exhibits intricate spatiotemporal dependencies and distributional complexities, which require generative models to effectively capture and model fine-grained spatial correlations and rapid temporal variations in network dynamics.
@@
Unfortunately, the aforementioned research can only generate coarse-grained mobile traffic with the same single spatiotemporal resolution as the training set. When generating and updating multi-scale traffic data, different models must be used to simulate different network layers, which greatly increases computational overhead. Moreover, generating the complex spatiotemporal distribution of fine-grained traffic exceeds the capabilities of existing methods. It requires advanced user localization techniques [10.1007/978-3-030-33607-3_45] or specialized network probes [BOZ2020107158], incurring substantial communication overhead and computational costs [9155340, GUERRERO201011], making it infeasible for continuous collection in large-scale live networks. A promising yet largely unexplored alternative is to design a unified model capable of jointly generating multi-scale mobile traffic, allowing the easily estimated coarse spatiotemporal granularity traffic to progressively guide the generation of finer-grained traffic, thereby achieving multi-scale traffic generation while improving the accuracy of fine-grained traffic synthesis.
@@
Recently, the emergence of diffusion models has offered promising insights for fine-grained traffic generation. Diffusion models operate by first progressively injecting noise into the original data and subsequently denoising step-by-step to reconstruct high-resolution images from pure noise. This denoising paradigm naturally aligns with the goal of fine-grained mobile traffic synthesis. Inspired by this, we aim to explore a control mechanism for the intermediate states in the noise-adding and denoising processes, as shown in Fig. 2, aligning the variation of noise intensity with the change in spatiotemporal resolution of mobile traffic. We seek to make the denoising process progressively zooms in on the mobile traffic, revealing increasingly finer spatiotemporal details at each step.
@@
Figure 2: Align the spatiotemporal refinement relationships of multi-scale mobile traffic with the denoising process of diffusion models.
@@
However, achieving this goal presents two main challenges. First, it remains unclear how to design effective denoising procedures within the diffusion framework to accurately represent multi-scale mobile traffic. Second, it is challenging to model and preserve the cross-resolution dependencies and statistical constraints between traffic at different granularities.
@@
To address the above challenges, we propose ZoomDiff, a multi-resolution mobile traffic generation model. It achieves a mapping from Urban Environmental Context (UEC) to multi-level mobile traffic, allowing for a zoom-like progressive generation for traffic with different scales. First, we design Denoising Refinement Diffusion Models (DRDM). It is a multi-stage diffusion framework, enabling controllable denoising intermediates through a Hierarchical Noise-adding Process (HNaP) and a Resolution Refinement Denoising Process (RRDP). Multi-scale traffic with different resolutions can be directly extracted from intermediate states at corresponding diffusion steps. Second, we design a diffusion prior noise mechanism to establish correlations between adjacent noise-adding and denoising stages. This design enables the coarse-resolution traffic to serve as guidance to control the generation of high-resolution traffic. Moreover, we discussed the noise scheduling strategies and the Resolution Greediness Preference of the multi-stage model required for implementing DRDM, and revised the corresponding strategies in conventional diffusion models. We found that the combination of discrete noise intensity, discrete noise-adding, and discrete denoising strategies enables ZoomDiff to achieve optimal multi-scale generation performance. The main contributions can be summarized as follows:
@@
  • To the best of our knowledge, we propose the first unified model capable of simultaneous generation of multi-scale mobile traffic. Leveraging the sophisticated design of DRDM, it aligns the progressive refinement of spatiotemporal scales of mobile traffic with the denoising process. In other words, multiple intermediate states of the denoising process are constrained to correspond to multi-scale mobile traffic.
@@
This paper is organized as follows. Section II provides an overview of related works on mobile traffic generation and spatiotemporal series prediction. Section III illustrates the preliminaries of our work, including an observation of the traffic correlations, the problem formulation, and an introduction to canonical diffusion models. In Section IV, we propose our ZoomDiff models and give detailed technical descriptions of the Environment-aware Conditional Information and the designed DRDM. Section V shows the evaluation results to verify ZoomDiff’s multi-scale generation accuracy and generalizability. Section VI discusses ZoomDiff’s capabilities in BS traffic refinement and cross-city generalization performance. Section VII provides a conclusion for this paper.
@@ II Related Work
II-A Mobile Traffic Generation
@@
Mobile traffic generation aims to learn the environment-aware spatiotemporal dynamics by mining the correlation between mobile traffic and external conditions. Traditional generative algorithms, including mathematical statistics [10.1145/1159913.1159928, 10.1145/1140086.1140094] and simulation-based approaches [10.1145/1028788.1028798, 10.1145/1081081.1081129, 10.1145/1231956.1231970], have limited generalization capabilities. With the development of Machine Learning (ML), generative models such as Generative Adversarial Networks (GANs) [10.1145/3485983.3494844, 5GT-GAN, Hierarchical-GAN, keGAN] and Transformers [10946853, UniST] can capture the complex correlations between network demand and environmental features in a data-driven manner. Pandey et al. [5GT-GAN] and Li et al. [Hierarchical-GAN] employed GANs to autoregressively model the temporal dynamics of traffic sequences for synthetic traffic generation. Hui et al. [keGAN] integrated urban environments into Urban Knowledge Graphs (UKGs) for generation guidance. As the stable generative capabilities of diffusion models are validated, recent works have attempted to replace GANs with diffusion-based generators. Zhou et al. [KSTDiff] applied UKG to the diffusion denoising process for accurate urban population generation. Sivaroopan et al. [NetDiffus] and Jiang et al. [NetDiffusion] converted one-dimensional mobile traffic into two-dimensional images to leverage the strong image generation capabilities of diffusion models.
@@
However, existing research primarily focuses on traffic generation at a single spatiotemporal resolution. For the task of generating multi-scale traffic, multiple models or separate generations are typically required. In addition, these models are designed for coarse-scale traffic, such as BS or cell levels, and perform poorly when generating fine-grained traffic at small grid scales. In contrast, ZoomDiff can generate multi-scale mobile traffic in a single process and progressively guide the generation of fine-grained traffic through coarse-scale traffic, achieving both improved generation accuracy and enhanced efficiency.
II-B Spatiotemporal Series Prediction/Generation
Mobile traffic generation can be regarded as a controllable spatiotemporal generation task. Numerous studies have employed models such as Convolutional Neural Network (CNN) [CNN-1, CNN-2, CNN-3], Recurrent Neural Network (RNN) [RNN-1, RNN-2, RNN-3], and Transformers [Tran-1, Trans-2, Trans-3, Trans-4] for spatiotemporal prediction and generation. Based on GANs, Yoon et al. [TimeGAN] proposed TimeGAN, which combines the flexibility of unsupervised paradigms with the control of supervised training by jointly optimizing supervision and adversarial learning in the embedding space. Saxena et al. [2019arXiv190708556S] introduced D-GAN, which jointly learns data generation and variational inference. In recent years, an increasing number of works have explored diffusion models to capture spatiotemporal distribution patterns. Yuan et al. [10.1145/3580305.3599511] improved the spatiotemporal point process (STPP) using diffusion models to learn complex joint spatiotemporal distributions in urban data. Yuan et al. [Diffusion-TS] reconstructed samples at each diffusion step and introduced Fourier-based loss terms to capture temporal dynamics. In addition to direct modeling in the temporal and spatial domains, learning spatiotemporal patterns through state space is becoming increasingly popular. Inspired by deep state space models (S4), Zhou et al. [10.5555/3618408.3620205] proposed LS4, which enhances spatiotemporal modeling capabilities via latex Ordinary Differential Equation (ODE) evolution in state space.
Researchers have attempted to achieve multi-resolution prediction of time series. Shen et al. [ICLR2024_d64740dd] utilized cascaded diffusion models that progressively refine the temporal resolution of the series. Shabani et al. [2022arXiv220604038S] proposed Scaleformer, in which a set of Transformer modules is used to iteratively refine time series across different temporal scales. However, these cascaded architectures linearly increase the number of model parameters as the resolution hierarchy grows and are prone to error accumulation. Fan et al. [2024arXiv240305751F] proposed MG-TSD, which leverages the inherent temporal granularity levels within the data as predefined targets at intermediate diffusion steps. Nevertheless, this model focuses solely on generating time-domain multi-resolution sequences without incorporating spatial multi-scale information. Li et al. [MDPNet] introduced MDPNet, which leverages spatial hierarchy to guide an ODE system in learning spatial semantics and performs forecasting using a diffusion-based encoder. However, the multi-scale spatial information in this approach only serves as auxiliary input for understanding spatial distributions, and the model does not produce multi-resolution outputs.
@@ III PRELIMINARY
III-A Correlations among Multi-scale mobile traffic
The mobile traffic across different scales, including the BS traffic, cell traffic, and grid-level traffic, exhibits a top-down aggregation relationship due to the inherent hierarchy among their service areas. Figure 4 shows the temporal decomposition of hourly BS traffic, cell traffic, and 50m grid traffic from three kinds of Area of Interest (AoI) over one week. The trend component reflects the distribution of daily average traffic over seven days, while the seasonal component captures the temporal patterns of daily traffic after removing the trend. To be quantitative, Fig. 3 computes the RV-coefficients [10.2307/2347233] between adjacent levels of mobile traffic. It can be observed that traffic at neighboring scales within the same AOI exhibits high similarity. This provides the theoretical foundation for jointly generating multi-scale traffic, where coarse-grained traffic can guide the generation of fine-grained traffic. The RV-coefficient is defined as:
@@ III-B Problem Formulation
Mobile traffic at different layers of network nodes reflects user demand distributions at different spatiotemporal scales. Given the UEC $\boldsymbol{c}$, the objective of the model is to learn a mapping from UEC to the corresponding regional mobile traffic $\boldsymbol{x}$, and simultaneously generate multi-level traffic within the region at different spatiotemporal resolutions. Consider a mobile network with $K\geq 2$ levels, let $\boldsymbol{x}_{s_{\delta_{k}}t_{\tau_{k}}}$ denote the mobile traffic at level $k$ with spatial resolution $s_{\delta_{k}}$ and temporal resolution $t_{\tau_{k}}$, where $1\leq k\leq K$, $\delta_{k}\geq\delta_{k-1}$ and $\tau_{k}\geq\tau_{k-1}$. This mapping can be expressed as $\boldsymbol{c}\mapsto(\boldsymbol{x}_{s_{\delta_{1}}t_{\tau_{1}}},...,\boldsymbol{x}_{s_{\delta_{K-1}}t_{\tau_{K-1}}},\boldsymbol{x}_{s_{\delta_{K}}t_{\tau_{K}}})^{T\times H\times W}$, where $T$ denotes the temporal length of the finest traffic series, and $H$, $W$ represent the finest spatial dimensions. As the spatiotemporal resolution increases, the traffic becomes increasingly influenced by fine-grained urban environments and user behavioral characteristics. Therefore, generating high-dimensional spatiotemporal traffic at fine resolutions is a complex task.